Scientists are studying the temperature on a distant planet. They find that the surface temperature at one location is 35°
celsius. They also find that the temperature decreases by 3° celsius for each kilometer you go up from the surface. Let T represent the temperature (in Celsius), and let
H be the height above the surface (in kilometers). Write an equation relating to T to H, and then graph

Answer by Theo(13342)

the formula would be T = 35 - 3 * H.
when H = 0 (at the surface), the temperaturs is 35 degrees celsius.
for every 1 kilometer above the surface, the temperature goes down 3 degrees celsius.
when H = 0, the temperature is 35 degrees celsius.
when H = 1, the temperature is 32 degrees celsius.
when H = 2, the temperature is 29 degrees celsius.
H is the number of kilometers above the surface at the specified location.
T is the temperature.

the following graph shows T for h = 0 to 5.
in the graph, y takes the place of T and x takes the place of H.
same formula, just y = 35 - 3 * x rather than T = 35 - 3 * H.

$$$